# Anisotropic Green Coordinates

### v1

@@ 3.2. Construction of Anisotropic Green Coordinates
Before deriving the anisotropic Green coordinates, we first revisit the construction of the classical isotropic Green coordinates, which are defined on an oriented simplicial surface cage $P=(\mathbb{V},\mathbb{T})$ (2008Green; 2009DriveGreen), where $\mathbb{V}$ denotes the cage vertices and $\mathbb{T}$ represents the face normals. Green coordinates are constructed based on Green’s third identity. Specifically, when $\Omega$ is a bounded open set in $\mathbb{R}^{d}$ and $u(\eta)$ is a harmonic function on $\overline{\Omega}$, then for $\eta\in\Omega$, we have:
(9) $u(\eta)=\int_{\partial\Omega}{u(\xi)\frac{\partial G}{\partial\mathbf{n}_{\xi}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}-\int_{\partial\Omega}{G(\xi,\eta)\frac{\partial u}{\partial\mathbf{n}_{\xi}}(\xi)\ \mathrm{d}\sigma_{\xi}},$
where $G(\xi,\eta)$ is the fundamental solution of the isotropic Laplacian equation (Eq. (7)) with $\mathbf{u}=\xi-\eta$.
Remark: The left hand side of Eq. (9) can be written as $c(\eta)u(\eta)$ if $\eta\in\mathbb{R}^{d}$ is not constrained in the open set $\Omega$. For a smooth boundary $\Omega$, the coefficient $c(\eta)$ satisfies $c(\eta)=1$ if $\eta\in\Omega$, $c(\eta)=1/2$ if $\eta\in\partial\Omega$, and $c(\eta)=0$ if $\eta\in\mathbb{R}^{d}\setminus\overline{\Omega}$.
Given the cage $P=(\mathbb{V},\mathbb{T})$, we can use above identity to reproduce any point $\eta$ inside the cage using the vertices $\{\mathbf{v}_{i}|i\in I_{\mathbb{V}}\}$ and normals $\{\mathbf{n}(t_{j})|j\in I_{\mathbb{T}}\}$ of the cage boundary as follows:
(10) $\eta=F(\eta,P)=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)\mathbf{n}(t_{j})},\ \ \eta\in\Omega,$
@@
(11) $\phi_{i}(\eta)$ $=\int_{\xi\in\mathrm{N}\{\mathbf{v}_{i}\}}{\Gamma_{\mathbf{v}_{i}}(\xi)\frac{\partial G(\xi,\eta)}{\partial\mathbf{n}(\xi)}\ \mathrm{d}\sigma_{\xi},i\in I_{\mathbb{V}}},$
@@
Figure 1. Illustration of the hat function.
Here, $\mathrm{N}\{\mathbf{v}_{i}\}$ represents the faces adjacent to $\mathbf{v}_{i}$, and $\Gamma_{\mathbf{v}_{i}}(\xi)$ is a piecewise linear hat function, which takes the value one at the vertex $\mathbf{v}_{i}$, zero at other vertices, and varies linearly on each face adjacent to $\mathbf{v}_{i}$. As shown in Fig. 1, in the 2D case, $\Gamma_{\mathbf{v}_{i}}(\xi)$ is given by $\Gamma_{\mathbf{v}_{i}}(\xi)=\frac{\|\xi-\mathbf{v}_{i+1}\|}{\|\mathbf{v}_{i+1}-\mathbf{v}_{i}\|}$. In the 3D scenario, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ is expressed as $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)=1-\frac{(\xi-\mathbf{v}_{j_{1}})\cdot\mathbf{h}_{j_{1}}}{\|\mathbf{h}_{j_{1}}\|^{2}}$, where $\mathbf{h}_{j_{1}}$ is the height of the opposite edge of $\mathbf{v}_{j_{1}}$. Moreover, based on the area ratio, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ essentially represents the barycentric coordinate of $\xi$ relative to $\mathbf{v}_{j_{1}}$ within the triangle $\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$. Therefore, $\xi=\Gamma_{\mathbf{v}_{j_{1}}}(\xi)\mathbf{v}_{j_{1}}+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)\mathbf{v}_{j_{2}}+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)\mathbf{v}_{j_{3}}$ and $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)=1.0$.
When the source cage $P$ is deformed to the target cage $\tilde{P}$, the new location $\tilde{\eta}$ toward $\eta$ is given by
(13) $\tilde{\eta}=F(\eta,\tilde{P})=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)s_{j}\tilde{\mathbf{n}}(t_{j})},$
where $\tilde{\mathbf{v}}_{i}$ and $\tilde{\mathbf{n}}(t_{j})$ are the vertices and normals of the deformed cage, respectively. In the following context, the tilde symbol will consistently be used to represent quantities associated with the deformed value or target cage. $s_{j}$ is the scale factor to keep scale invariance in 2D and quasiconformality in 3D, which is detailed in Section 3 of  (2008Green).
Next, we derive the anisotropic boundary integral formulation of the anisotropic Laplacian equation, which establishes the mathematical foundation for anisotropic Green coordinates.
@@
If $\Omega$ is a bounded open set in $\mathbb{R}^{d}$, and $u(\xi)$ is a scalar function satisfying $\nabla\cdot(\mathbf{A}\nabla u(\mathbf{\xi}))=0$ on $\overline{\Omega}$, where $\mathbf{A}$ is symmetric positive definite. Then, for $\eta\in\Omega$,
@@
where $G_{\mathbf{A}}(\xi,\eta)$ is the fundamental solution of anisotropic Laplacian equation $\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta))=\delta(\xi-\eta)$.
@@
We begin the proof by applying the divergence theorem. Let $\Omega$ be a bounded open set in $\mathbb{R}^{d}$ and $\mathbf{F}$ is a smooth vector-valued function defined on the closure $\overline{\Omega}$. Then,
@@
We can derive the anisotropic Green coordinates with the natural linear reproduction property by utilizing Eq. (14). By setting $u(\eta)=\eta$, we can express $\eta\in\Omega$ in terms of the anisotropic boundary integration. There is a remark that $u$ is a scalar function in Eq. (14). However, we can separately apply the identity to each component along the coordinate axes. Specifically, let $u_{k}(\mathbf{\eta})=\eta_{k}$, which returns the $k$-th component of $\eta=(\eta_{1},\eta_{2},...,\eta_{d})^{\top}$. Moreover, let $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d})$, where $\mathbf{a}_{k}$ is the $k$-th column of $\mathbf{A}$. Then, $\mathbf{A}\nabla u_{k}(\xi)\cdot\mathbf{n}(\xi)=(\mathbf{a}_{k})^{\top}\mathbf{n}(\xi)$. We can then combine the $d$ components to obtain $\eta=(u_{1}(\eta),...,u_{d}(\eta))^{\top}$. Moreover, the composed normal term on the right-hand side is $\mathbf{A}^{\top}\mathbf{n}(t_{j})=\mathbf{A}\mathbf{n}(t_{j})$. Considering the boundary integration region as the cage boundary, we have the following equations:
@@
$-\sum_{j\in I_{\mathbb{T}}}{\mathbf{A}\mathbf{n}(t_{j})}\Big(\int_{t_{j}}{G_{\mathbf{A}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}\Big),$
where $I_{\mathbb{T}}$ represents the index set of the cage faces. By rearranging the above equation, the anisotropic Green coordinates can be written as:
(23) $\eta=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)\ \mathbf{A}\mathbf{n}(t_{j})},\eta\in\Omega,$
@@
(24) $\phi_{i}^{\mathbf{A}}(\eta)$ $=\int_{\xi\in\mathrm{N}\{\mathbf{v}_{i}\}}{\Gamma_{i}(\xi)\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi)\ \mathrm{d}\sigma_{\xi},\ i\in I_{\mathbb{V}}},$
@@
Compared to the traditional Green coordinates, the integral expressions of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are now related to $G_{\mathbf{A}}(\xi,\eta)$ instead of $G(\xi,\eta)$. Moreover, the Neumann term of Eq. (23), which is constructed based on the cage normals, now becomes $\mathbf{A}\mathbf{n}(t_{j})$. When the source cage is deformed to $\tilde{P}$ with new vertices $\tilde{\mathbf{v}}_{i}$ and face normals $\tilde{\mathbf{n}}(t_{j})$, the deformed position of $\eta$ is
(26) $\tilde{\eta}=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)s_{j}\mathbf{A}\tilde{\mathbf{n}}(t_{j})},$
where $s_{j}$ is the scale factor analogous to that in (2008Green). A detailed discussion of this term will be provided in the following subsection.
@@ 3.3. Properties
To analyze the deformation properties of the anisotropic Green coordinates (Eq. (26)), we first revisit the properties of the isotropic case, as examined in (2008Green; 2009DriveGreen). Specifically, these properties include: (1) Linear reproduction: $\eta=F(P,\eta)$ for $\eta\in\Omega$; (2) Translation invariance: $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)=1}$ for $\eta\in\Omega$; (3) Rotation and scale invariance: for an affine transformation $T$, $T\eta=F(\eta,TP)$; (4) Shape preservation: the deformation is angle-preserving in 2D and quasi-conformal in 3D; and (5) Smoothness: $\phi_{i}(\eta)$ and $\psi_{j}(\eta)$ are $C^{\infty}$ for $\eta\in\Omega$. We observe that the anisotropic Green coordinates retain the properties of linear reproduction, translation invariance, and scale invariance. However, when $\mathbf{A}\neq\mathbf{I}$, it does not satisfy the rotation invariance. Additionally, they are not angle-preserving in 2D.
Linear reproduction: The linear reproduction property is directly achieved through the anisotropic boundary integral formula. In Eq. (26), when the source and target cages are identical and $s_{j}=1$, substituting the coordinate expressions back yields Eq. (22), thereby confirming that $\tilde{\eta}=\eta$.
Translation invariance: Translation invariance is obtained from the fact that $\phi_{i}^{\mathbf{A}}(\eta)$ satisfies the partition of unity, i.e., $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)=1}$. This can be proven by setting $u(\eta)=1$ in Eq. (14). Then, we get
@@
Additionally, $\Gamma_{i}(\eta)$ in $\phi_{i}^{\mathbf{A}}(\eta)$ is actually the barycentric coordinates of the simplices associated with vertices $\mathbf{v}_{i}$. Therefore,
(28) $\sum_{i\in I_{\mathbb{V}}}\phi_{i}^{\mathbf{A}}(\eta)=$ $\sum_{j\in I_{\mathbb{T}}}\sum_{\mathbf{v}_{i}\in\mathbb{V}(t_{j})}\Big({\int_{t_{j}}{\Gamma_{\mathbf{v}_{i}}(\xi)\ (\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}}}$
@@
(29) $s_{j}=\frac{\sqrt{\|\tilde{u}\|^{2}\|v\|^{2}-2(\tilde{u}\cdot\tilde{v})+\|\tilde{v}\|^{2}\|u\|^{2}}}{\sqrt{8}\mathrm{area}(t_{j})},$
where $u,v$ and $\tilde{u},\tilde{v}$ are two pairs of the corresponding vectors of the triangle $t_{j}$ in the source and target cages, respectively.
Smoothness and generalized harmonic We also have the property that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are smooth for $\eta\in\Omega$. The reason is that $G_{\mathbf{A}}(\xi,\eta)$ is an elementary function with respect to $\eta$ and has no singularities within the integration domain, given that $\xi\neq\eta$. Consequently, $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are also smooth with respect to $\eta$, as they are obtained through differentiation and integration of $G_{\mathbf{A}}(\xi,\eta)$ (Note that $\Gamma_{i}(\xi)$ is independent of $\eta$). We also have $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\phi_{i}^{\mathbf{A}}(\eta))=0$ and $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\psi_{j}^{\mathbf{A}}(\eta))=0$. We call this property generalized harmonic with respect to $\mathbf{A}$. This can be briefly proven as follows: Due to the smoothness of $G_{\mathbf{A}}(\xi,\eta)$ when $\xi\neq\eta$, differentiation with respect to $\eta$ can be interchanged with the integration with respect to $\xi$. Moreover, the gradient of $G_{\mathbf{A}}(\xi,\eta)$ with respect to $\eta$ is simply the negative of its gradient with respect to $\xi$. Given that $G_{\mathbf{A}}(\xi,\eta)$ is generalized harmonic with respect to $\mathbf{A}$ in $\xi$, it follows that it is also generalized harmonic in $\eta$ in $\Omega$. It follows that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ also inherit this property.

### v2

@@ 3.2. Construction of Anisotropic Green Coordinates
Before deriving the anisotropic Green coordinates, we first revisit the construction of the classical isotropic Green coordinates, which are defined on an oriented simplicial surface cage $P=(\mathbb{V},\mathbb{T})$ (Lipman et al., 2008; Lipman and Levin, 2009), where $\mathbb{V}$ denotes the cage vertices and $\mathbb{T}$ represents the face normals. Green coordinates are constructed based on Green’s third identity. Specifically, when $\Omega$ is a bounded open set in $\mathbb{R}^{d}$ with $\mathcal{C}^{2}$ boundary, and $u(\eta)$ is a harmonic function on $\overline{\Omega}$, then for $\eta\in\Omega$, we have:
(9) $u(\eta)=\int_{\partial\Omega}{u(\xi)\frac{\partial G}{\partial\mathbf{n}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}-\int_{\partial\Omega}{G(\xi,\eta)\frac{\partial u}{\partial\mathbf{n}}(\xi)\ \mathrm{d}\sigma_{\xi}},$
where $G(\xi,\eta)$ is the fundamental solution of the isotropic Laplacian equation (7). The directional derivatives are taken along the outward normal.
Remark: The left hand side of Eq. (9) can be written as $c(\eta)u(\eta)$ if $\eta\in\mathbb{R}^{d}$ is not constrained in the open set $\Omega$ and $\partial\Omega$ is $\mathcal{C}^{2}$. The coefficient $c(\eta)$ satisfies $c(\eta)=1$ if $\eta\in\Omega$, $c(\eta)=1/2$ if $\eta\in\partial\Omega$, and $c(\eta)=0$ if $\eta\in\mathbb{R}^{d}\setminus\overline{\Omega}$. We ignore the boundary continuity condition for discrete inputs when $\eta\in\Omega$.
Figure 1. Illustration of the hat function in 2D and 3D.
Given the cage $P=(\mathbb{V},\mathbb{T})$, we can use above identity to reproduce any point $\eta$ inside the cage using the vertices $\{\mathbf{v}_{i}|i\in I_{\mathbb{V}}\}$ and outward normals $\{\mathbf{n}_{j}|j\in I_{\mathbb{T}}\}$ of the cage boundary as follows:
(10) $\eta=F(\eta,P)=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)\mathbf{n}_{j}},\ \ \eta\in\Omega,$
@@
(11) $\phi_{i}(\eta)$ $=\int_{\xi\in\mathcal{N}\{\mathbf{v}_{i}\}}{\Gamma_{\mathbf{v}_{i}}(\xi)\frac{\partial G}{\partial\mathbf{n}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi},i\in I_{\mathbb{V}}},$
@@
In the above formulations, $I_{\mathbb{V}}$ and $I_{\mathbb{T}}$ represent the index set of the cage vertices and faces, respectively; $\mathcal{N}\{\mathbf{v}_{i}\}$ represents the faces adjacent to $\mathbf{v}_{i}$, and $\Gamma_{\mathbf{v}_{i}}(\xi)$ is a piecewise linear hat function, which takes the value one at $\mathbf{v}_{i}$, zero at all other vertices, and varies linearly across each face within $\mathcal{N}\{\mathbf{v}_{i}\}$. As illustrated in Fig. 1, the explicit form of $\Gamma_{\mathbf{v}_{i}}(\xi)$ in the 2D case is given by $\Gamma_{\mathbf{v}_{i}}(\xi)=\frac{\|\xi-\mathbf{v}_{i+1}\|}{\|\mathbf{v}_{i+1}-\mathbf{v}_{i}\|}$. In the 3D scenario, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ is expressed as $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)=1-\frac{(\xi-\mathbf{v}_{j_{1}})\cdot\mathbf{h}_{j_{1}}}{\|\mathbf{h}_{j_{1}}\|^{2}}$, where $\mathbf{h}_{j_{1}}$ denotes the altitude corresponding to the edge opposite $\mathbf{v}_{j_{1}}$. Furthermore, $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)$ essentially represents the barycentric coordinate of $\xi$ relative to $\mathbf{v}_{j_{1}}$ within $\triangle\mathbf{v}_{j_{1}}\mathbf{v}_{j_{2}}\mathbf{v}_{j_{3}}$. That is, $\xi=\Gamma_{\mathbf{v}_{j_{1}}}(\xi)\mathbf{v}_{j_{1}}+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)\mathbf{v}_{j_{2}}+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)\mathbf{v}_{j_{3}}$ and $\Gamma_{\mathbf{v}_{j_{1}}}(\xi)+\Gamma_{\mathbf{v}_{j_{2}}}(\xi)+\Gamma_{\mathbf{v}_{j_{3}}}(\xi)=1.0$.
When the source cage $P$ is deformed to the target cage $\tilde{P}$, the new location $\tilde{\eta}$ of $\eta$ is computed as:
(13) $\tilde{\eta}=F(\eta,\tilde{P})=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}(\eta)s_{j}\tilde{\mathbf{n}}_{j}},$
where $\tilde{\mathbf{v}}_{i}$ and $\tilde{\mathbf{n}}_{j}$ denote the vertices and normals of the deformed cage, respectively. Throughout this work, the tilde symbol is used to represent quantities associated with the deformed value or target cage. The scale factor $s_{j}$ preserves scale invariance in 2D and quasiconformality in 3D, as detailed in Section 3 of  (Lipman et al., 2008).
Next, we derive the anisotropic boundary integral formulation of the anisotropic Laplacian equation, which establishes the mathematical foundation for our anisotropic Green coordinates.
@@
If $\Omega$ is a bounded open set in $\mathbb{R}^{d}$ with $\mathcal{C}^{2}$ boundary, and $u(\xi)$ is a $\mathcal{C}^{2}$ scalar function satisfying $\nabla\cdot(\mathbf{A}\nabla u(\mathbf{\xi}))=0$ on $\overline{\Omega}$, where $\mathbf{A}$ is symmetric positive definite. Then, for $\eta\in\Omega$, we have
@@
where $G_{\mathbf{A}}(\xi,\eta)$ is the fundamental solution of the anisotropic Laplacian equation $\nabla_{\xi}\cdot(\mathbf{A}\nabla_{\xi}G_{\mathbf{A}}(\xi,\eta))=\delta(\xi-\eta)$, and $\mathbf{n}(\xi)$ the outward normal of $\partial\Omega$ at $\xi$.
@@
We begin the proof by applying the divergence theorem. Let $\Omega$ be a bounded open set in $\mathbb{R}^{d}$ and $\mathbf{F}$ is a $\mathcal{C}^{2}$ vector-valued function defined on the closure $\overline{\Omega}$. Then,
@@
Remark: From a strict mathematical perspective, the proof above remains not rigorous enough, as $G_{\mathbf{A}}$ is singular at $\xi=\eta$ when performing the volume integration. The standard approach is to excise $B_{\epsilon}(\eta)$ centered at $\eta$ and let $\varepsilon\rightarrow 0$, or to construct a sequence of functions that converges pointwise to the fundamental solution while remaining free of singularities (replacing $\|\xi-\eta\|$ with $(\|\xi-\eta\|^{2}+\varepsilon^{2})^{1/2}$. Here, however, we prove the result directly in the sense of distributions.
We can derive the anisotropic Green coordinates with the natural linear reproduction property by setting $u(\eta)=\eta$ in Eq. (14). There is a remark that $u$ is a scalar function in Eq. (14). However, we can apply the identity to each coordinate component separately. Specifically, let $u_{k}(\mathbf{\eta})=\eta_{k}$, which returns the $k$-th component of $\eta=(\eta_{1},\eta_{2},...,\eta_{d})^{\top}$. Moreover, denote $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},...,\mathbf{a}_{d})$, where $\mathbf{a}_{k}$ is the $k$-th column of $\mathbf{A}$. Then, $\mathbf{A}\nabla u_{k}(\xi)\cdot\mathbf{n}(\xi)=(\mathbf{a}_{k})^{\top}\mathbf{n}(\xi)$. We can then combine the $d$ components to obtain $\eta=(u_{1}(\eta),...,u_{d}(\eta))^{\top}$. Moreover, the composed normal term on the right-hand side is $\mathbf{A}^{\top}\mathbf{n}_{j}=\mathbf{A}\mathbf{n}_{j}$. Considering the boundary integration region as the cage boundary, we have the following equations:
@@
$-\sum_{j\in I_{\mathbb{T}}}{\mathbf{A}\mathbf{n}_{j}}\Big(\int_{t_{j}}{G_{\mathbf{A}}(\xi,\eta)\ \mathrm{d}\sigma_{\xi}}\Big),$
where $I_{\mathbb{T}}$ denotes the index set of the cage faces, and $\mathbb{V}(t_{j})$ represents the vertex set of $t_{j}$. By rearranging the above equation, the anisotropic Green coordinates can be expressed as:
(23) $\eta=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\mathbf{v}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)\ \mathbf{A}\mathbf{n}_{j}},\eta\in\Omega,$
@@
(24) $\phi_{i}^{\mathbf{A}}(\eta)$ $=\int_{\xi\in\mathcal{N}\{\mathbf{v}_{i}\}}{\Gamma_{i}(\xi)\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi)\ \mathrm{d}\sigma_{\xi},\ i\in I_{\mathbb{V}}},$
@@
Compared to the traditional Green coordinates (Eqs (10) to (12)), the integral expressions of $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are now related to $G_{\mathbf{A}}(\xi,\eta)$ instead of $G(\xi,\eta)$. Moreover, the Neumann term of Eq. (23), which is constructed based on the cage normals, now becomes $\mathbf{A}\mathbf{n}_{j}$. When the source cage is deformed to $\tilde{P}$ with new vertices $\tilde{\mathbf{v}}_{i}$ and face normals $\tilde{\mathbf{n}}_{j}$, the deformed position $\tilde{\eta}$ of $\eta$ is given by:
(26) $\tilde{\eta}=\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)\tilde{\mathbf{v}}_{i}}+\sum_{j\in I_{\mathbb{T}}}{\psi_{j}^{\mathbf{A}}(\eta)s_{j}\mathbf{A}\tilde{\mathbf{n}}_{j}},$
where $s_{j}$ is the scale factor analogous to that in (Lipman et al., 2008). A detailed discussion of this term will be provided in the following subsection.
@@ 3.3. Properties
To analyze the deformation properties of the anisotropic Green coordinates, we first revisit the properties of the isotropic case, as examined in (Lipman et al., 2008; Lipman and Levin, 2009). Specifically, these properties include: (1) Linear reproduction: $\eta=F(P,\eta)$ for $\eta\in\Omega$; (2) Translation invariance: $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}(\eta)=1}$ for $\eta\in\Omega$; (3) Rotation and scale invariance: for an affine transformation $T$, $T\eta=F(\eta,TP)$; (4) Shape preservation: the deformation is angle-preserving in 2D and quasi-conformal in 3D; and (5) Smoothness: $\phi_{i}(\eta)$ and $\psi_{j}(\eta)$ are $C^{\infty}$ for $\eta\in\Omega$. We observe that the anisotropic Green coordinates retain the properties of linear reproduction, translation invariance, scale invariance, quasi-conformality and smoothness. However, it does not satisfy the rotation invariance. Additionally, in 2D, these coordinates are not angle-preserving but exhibit quasi-conformality.
Linear reproduction: The anisotropic boundary integral formula naturally satisfies the linear reproduction property. As shown in Eq. (26), when the source and target cages are identical and $s_{j}=1$, substituting the coordinate expressions back yields Eq. (22), verifying that $\tilde{\eta}=\eta$.
Translation invariance: Translation invariance follows from the fact that $\phi_{i}^{\mathbf{A}}(\eta)$ satisfies the partition of unity property, i.e., $\sum_{i\in I_{\mathbb{V}}}{\phi_{i}^{\mathbf{A}}(\eta)=1}$. This property can be proven by setting $u(\eta)=1$ in Eq. (14), which yields:
@@
Furthermore, $\Gamma_{\mathbf{v}_{i},t_{j}}(\eta)$ in the expression for $\phi_{i}^{\mathbf{A}}(\eta)$ is actually the barycentric coordinates of $\mathbf{v}_{i}$ with respect to $t_{j}$. Therefore,
(28) $\sum_{i\in I_{\mathbb{V}}}\phi_{i}^{\mathbf{A}}(\eta)=$ $\sum_{j\in I_{\mathbb{T}}}\sum_{\mathbf{v}_{i}\in\mathbb{V}(t_{j})}\Big({\int_{t_{j}}{\Gamma_{\mathbf{v}_{i},t_{j}}(\xi)\ (\mathbf{A}\nabla G_{\mathbf{A}}(\xi,\eta)\cdot\mathbf{n}(\xi))\ \mathrm{d}\sigma_{\xi}\Big)}}$
@@
(29) $s_{j}=\sqrt{\frac{\sigma_{1}^{2}+\sigma_{2}^{2}}{2}}=\frac{\sqrt{\|\tilde{u}\|^{2}\|v\|^{2}-2(\tilde{u}\cdot\tilde{v})+\|\tilde{v}\|^{2}\|u\|^{2}}}{\sqrt{8}\mathrm{area}(t_{j})},$
where $\sigma_{1}$ and $\sigma_{2}$ denote the singular values of the linear map that transforms $t_{j}$ to $\tilde{t}_{j}$ when restricted to a plane. Moreover, $u,v$ and $\tilde{u},\tilde{v}$ represent the edge vectors of $t_{j}$ and $\tilde{t}_{j}$, respectively.
Quasi-conformality Although anisotropic Green coordinates are not strictly conformal in two dimensions, their quasi-conformality in both 2D and 3D can be established through the analysis of their relationship with the isotropic formulation presented in Section 5. Consequently, we also include the analysis of quasi-conformality in Section 5.
Smoothness and generalized harmonicity We also have the property that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are smooth for $\eta\in\Omega$. The reason is that $G_{\mathbf{A}}(\xi,\eta)$ is an elementary function with respect to $\eta$ and has no singularities within the integration domain as $\xi\neq\eta$. Consequently, $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ are also smooth with respect to $\eta$, as they are obtained through differentiation and integration of $G_{\mathbf{A}}(\xi,\eta)$ with respect to $\xi$, and $\Gamma_{i}(\xi)$ is independent of $\eta$. Moreover, we also conclude that $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\phi_{i}^{\mathbf{A}}(\eta))=0$ and $\nabla_{\eta}\cdot(\mathbf{A}\nabla_{\eta}\psi_{j}^{\mathbf{A}}(\eta))=0$. We call this property generalized harmonic with respect to $\mathbf{A}$. This can be briefly proven as follows: Due to the smoothness of $G_{\mathbf{A}}(\xi,\eta)$ when $\xi\neq\eta$, differentiation with respect to $\eta$ can be interchanged with the integration with respect to $\xi$. Moreover, the gradient of $G_{\mathbf{A}}(\xi,\eta)$ with respect to $\eta$ is simply the negative of its gradient with respect to $\xi$. Given that $G_{\mathbf{A}}(\xi,\eta)$ is generalized harmonic with respect to $\mathbf{A}$ in $\xi$, it is also generalized harmonic with respect to $\eta$ in $\Omega$. It follows that $\phi_{i}^{\mathbf{A}}(\eta)$ and $\psi_{j}^{\mathbf{A}}(\eta)$ also inherit this property.